OpenStudy (anonymous):

24/36 x 9/27 simplify? help plz

OpenStudy (anonymous):

24/36 * 9/27 = 2/3 * 1/3 = 2/9 or if you don't want to reduce at the beginning you can just multiply and reduce at the end 24/36 * 9/27 = 216/972 = 108/486 = 54/243 = 2/9

OpenStudy (anonymous):

thank u very much @mahnoorkeen2001

OpenStudy (anonymous):

welcome